# Mergers all the way down: stellar collisions and kinematics of a dense hierarchically forming massive star cluster in a dwarf starburst

### v1

@@ 3.1 Properties of young star clusters
Approximately 850 star clusters that fill our selection criteria form during the $\sim 7.5$ Myr of starburst. Star clusters above $\sim 300$ $\mathrm{M}_{\odot}$ exhibit a cluster mass function that has a power-law index between $-1.9$ and $-2$ throughout the starburst, fit using 20 bins with equal number of 40–50 clusters per bin. The slope of the mass function agrees with observed mass functions and luminosity functions across environments from low-metallicity galaxies such as the Magellanic Clouds to starburst systems (Elmegreen & Falgarone, 1996; Elmegreen & Efremov, 1997; Zhang & Fall, 1999; Hunter et al., 2003; Larsen, 2009; Randriamanakoto et al., 2013).
By the final snapshot, the majority of the clusters are already gas-free while all clusters have mean stellar ages less than 4.3 Myr. The most massive cluster at a mean stellar age of 1.9 Myr still has a gas-embedded core that is forming stars. While the tail end of the starburst is still ongoing, we limit the analysis only until this epoch. The majority of the stellar mass has already formed, and the main stellar collision product (Section 3.1.2) in the most massive cluster has not yet reached a mass significantly above 1000 $\mathrm{M}_{\odot}$, i.e. more than a factor of two beyond the upper stellar mass limit adopted in our VMS models (see Section 2.1.4). Fig. 1 shows the projected half-mass radii ($R_{\mathrm{eff}}$) and the corresponding surface densities $\Sigma_{\mathrm{eff}}=M_{\mathrm{cluster}}/(2\pi R_{\mathrm{eff}}^{2})$ in the $x$$y$ plane for the simulated gravitationally bound star clusters in the final snapshot. The results are compared to masses and half-light radii of young ($<10$ Myr) star clusters in the LEGUS survey (Brown & Gnedin, 2021) and of resolved or marginally resolved star clusters measured through gravitational lensing at redshifts $z\gtrsim 2.4$ (Vanzella et al. 2022b, a, 2023; Adamo et al. 2024; Mowla et al. 2024; Messa et al. 2025). The surface densities in the observed clusters have been computed using the reported half-light radii. We also show the overall best-fit mass-size relation from Brown & Gnedin (2021).
[…]
@@ 3.1.1 Stellar collisions and two TDEs
In total $\sim 12500$ stars more massive than $8$ $\mathrm{M}_{\odot}$ form throughout the starburst. 69 are more massive than $100$ $\mathrm{M}_{\odot}$. When stellar collisions occur in our simulation, only one star grows per cluster after segregating into the cluster core, similar to results in the literature (Portegies Zwart et al., 1999; Baumgardt & Klessen, 2011; Fujii & Portegies Zwart, 2013a; Rantala et al., 2024). In two cases the merger product of two stars below the limit of 100 $\mathrm{M}_{\odot}$ exceeds the VMS limit. The rest of the events that grow VMSs start with a star originally above 100 $\mathrm{M}_{\odot}$. The most massive star cluster hosts $>2000$ stars more massive than $8$ $\mathrm{M}_{\odot}$, up to $\sim 18$ of which are more massive than $100$ $\mathrm{M}_{\odot}$ and present in the cluster at the same time. 12 remain in the cluster in the final snapshot. 12 of the VMSs form after the main progenitor has reached 25% of its final mass and they form within $\sim 1$ pc of the cluster centre. The rest arrive via sub-cluster mergers. By the last snapshot, $78$ stars in the entire galaxy more massive than $40$ $\mathrm{M}_{\odot}$ have collapsed as black holes, out of which $8$ remain in the most massive cluster. 20 stars have exploded as pair-instability supernovae.
@@
We track the collisions of stars throughout the simulation for pairs that include a star or a black hole that had an initial stellar mass of at least $9$ $\mathrm{M}_{\odot}$. During the $\sim 7$ Myr of starburst there are in total $90$ collisions throughout the simulation, out of which $16$ involve two components both of which are more massive than $40$ $\mathrm{M}_{\odot}$. Almost $30\%$ of the collisions are between fairly equal mass objects (mass ratio $>1/3$). We note that compared to the hierarchical direct N-body simulations of Rantala et al. (2024), we do not see the growth of VMSs through collisions in the in-falling sub-clusters during the assembly of the most massive cluster. This may be due to somewhat delayed mass segregation and sub-cluster core collapse in our collisional dynamics implementation, compared to full direct N-body (see Fig. 1 in Lahén et al. 2025). The core collapse time scale, for instance, affects when VMS growth can occur in the sub-clusters (Fujii & Portegies Zwart, 2013b). Optionally the spread in stellar ages in each cluster due to the star formation history may delay the birth of VMSs and their expansion into the giant phase compared to the uniform age initial conditions adopted in N-body models. This shifts the collisional growth of VMSs to the later, more centrally concentrated stage of cluster assembly.
@@ 3.1.2 The most massive star
In the most massive cluster, one star grows to a peak mass of 1046 $\mathrm{M}_{\odot}$ through repeated collisions with other massive stars. It began its life as a 150 $\mathrm{M}_{\odot}$ star, and consequently merged with stars of 115 $\mathrm{M}_{\odot}$, 170 $\mathrm{M}_{\odot}$, 191 $\mathrm{M}_{\odot}$, 232 $\mathrm{M}_{\odot}$, 153 $\mathrm{M}_{\odot}$ and 44 $\mathrm{M}_{\odot}$ during a period of 1 Myr. The collisions are spaced by 0.1 Myr, 0.15 Myr, 0.4 Myr, 0.3 Myr and 4.4 kyr. The most massive star also merged with 13 low-mass stars ($<0.4$ $\mathrm{M}_{\odot}$), increasingly frequently during the final 0.2 Myr as it grew to the supergiant phase. We limit the analysis to this epoch, however the star would still have the potential to collide with a variety of other massive stars, including the most massive normal (i.e. not collisionally formed) VMS of 330 $\mathrm{M}_{\odot}$. This star is at a distance of only 0.05 pc from the cluster core and the most massive star in the final snapshot.
@@
The final VMS mass that far exceeds the upper stellar mass limit of the adopted IMF, and its growth being dominated by collisions with other VMSs, are consistent with hierarchical direct N-body models of star cluster assembly (Rantala et al., 2024). All of the individual collision partners are within the range of initial masses of the most massive stars in R136, therefore the physical scenario for the VMS growth through collisions is very well motivated based on observations. Why no such collisions are observed in R136 may be due to its too low central/effective density, or its higher metallicity (factor of tens) that allows stronger mass-loss and smaller stellar radii at evolved stages.
@@
Our resulting growth of the most massive star is less extreme than what is found in idealised cloud-scale hydrodynamical simulations (e.g. Reinoso et al. 2023; Fujii et al. 2024; Chon & Omukai 2024). Such simulations of dense star-forming clouds find VMSs in excess of $10^{3}$$10^{4}$ $\mathrm{M}_{\odot}$ forming through collisions and/or gas accretion. Future work with improved treatment for the evolution of the most massive stars will help address whether this difference is a result of the galactic environment and more complex cloud structure in our model, the stellar evolution models, the adopted IMF, or perhaps something else.
@@
Though we have refrained from making assumptions about the detailed evolution of stars beyond $500$ $\mathrm{M}_{\odot}$, an order of magnitude estimate for the final mass of the most massive star and its black hole remnant can be made based on simplified arguments about the mass and metallicity dependent wind mass loss e.g. by Vink (2018) and Sabhahit et al. (2023). The most massive star undergoes its last major collision at an age of $\sim 1.2$ Myr, and it can be expected to still live for one million years. A $\sim 1000$ $\mathrm{M}_{\odot}$ star at $\sim 0.01$ $\mathrm{Z}_{\odot}$ would initially lose mass in optically thick winds at a rate between $\sim 10^{-4.3}$$10^{-3.7}$ $\mathrm{M}_{\odot}$ yr-1 depending on the rate estimate. The star would lose in total between 25similar-toabsent25$\sim$ 25∼ 25160160160160 $\mathrm{M}_{\odot}$ during the final 1 Myr of its life, resulting in a black hole with a mass of at least $\sim 840$ $\mathrm{M}_{\odot}$. Even if the star would be rotating fast as a consequence of its merger history and therefore lose mass at a factor of 2–3 faster rate as argued by Gieles et al. (2025), it would still result in a final black hole mass of at least $640$ $\mathrm{M}_{\odot}$. Additionally, if the stellar collisions were implemented with a prescription for mass loss, we can approximate the star to end up with 10% lower maximum mass, $\sim 900$ $\mathrm{M}_{\odot}$, after the last major collision. This would further decrease the resulting black hole mass to a minimum of 600 $\mathrm{M}_{\odot}$ (rotating VMS with Sabhahit et al. 2023 mass loss rate) or 770 $\mathrm{M}_{\odot}$ (without rotation). The result would therefore be an intermediate mass black hole in a proto-globular cluster-like system. Whether such a black hole would be retained in the cluster over several Gyrs would require a more detailed modelling of black hole–black hole interactions and mergers in the simulation.
@@
Figure 3: The radial projected profiles of stellar mass surface density (a), V-band surface brightness (b), 3D mass density (c) and stellar number density (d) of the most massive cluster at the four epochs shown in Fig. 2 and in the last snapshot when the cluster has a mean stellar age of 1.9 Myr. The solid lines in panels a, c and d indicate the gravitationally bound stellar component and the dashed lines show the total stellar mass in the region. In panel b the solid and dashed lines show the attenuated and the unattenuated surface brightness profiles, respectively. The power-law slopes of $-0.5$, $-1$, $-2$ and $-3$ are indicated in the panels. The best-fit surface density and brightness profiles of R136 from McLaughlin & van der Marel (2005) are shown in black in panels a and b, and the right-side y-axis in panel b shows the surface brightness scale in mag arcsec-2. The vertical bars at the bottom of the panels show the projected half-mass radii of the bound component (panels a, c and d) and the projected half-light radii of the attenuated and unattenuated light profiles (b) with respective linestyles.
@@ 3.2 Properties of the most massive cluster
The stellar light of the main cluster is attenuated and the other massive clump in the images remains the brightest cluster in V-band for the duration of this coalescence, irrespective of the masking applied in the centring. The V-band attenuation $A_{V}$ is $\sim 0.9$ mag at these epochs, computed as the AB-magnitude difference between the unattenuated and attenuated images within a radius of 1 pc of the centre of light of the unattenuated image. This is a moderate value for observed star clusters and not as extreme as the most attenuated clusters observed in some dwarf galaxies (e.g. Calzetti et al. 2023). The reddening $E(B-V)$ within the same 1 pc region is 0.7 and 0.5 in the two first epochs, similar to what is typically found for young clusters in dwarf galaxies or starbursting systems ($E(B-V)$ up to 0.7 e.g. in Adamo et al. 2020; Whitmore et al. 2020).
The last two epochs in Fig. 2 show centrally concentrated star formation. The central 1 pc has a star-forming disky gas component until the last snapshot. The gas disk is being fed along a gaseous filament that extends toward the cluster from north-west in the rightmost panels. $A_{V}$ peaks at $\sim 1.4$ during the last epochs, and $E(B-V)$ drops to 0.3–0.4.
@@ 3.2.1 Density, surface density and surface brightness profiles
The density and brightness profiles of the most massive cluster are consistent with power-law shapes especially in the outer parts and at later epochs when the stellar distribution is less clumpy (see Fig. 2). The infalling clusters in our simulation can be seen as an increase in density at intermediate radii (11113333 pc) at earlier epochs. The outer radial power-law slope in surface density is $-2.2$, as expected from models of hierarchical sub-clusters assembly that predict slopes that approach $-2$ (Sills et al., 2018; Grudić et al., 2018; Rantala et al., 2024). The outer power-law has a respective density slope of slightly shallower than $-3$ and continues well into the sub-parsec regime. The final central 3D mass density, 3D number density and surface mass density (inner $\sim 1000$ stars) are a factor of $\sim 190$, 10 and 15 times higher than the half-mass values shown in Fig. 1, respectively. The unattenuated central surface brightness profile likewise has a steep inner slope, dominated by the most massive stars in the cluster core. Based on these simulated brightness and density profiles, we therefore infer that the central surface densities $\Sigma_{0}$ of the observed high-redshift clusters may be at least an order of magnitude higher than the effective values $\Sigma_{\mathrm{eff}}$, i.e. $\Sigma_{0}>10^{6}$ $\mathrm{M}_{\odot}$ pc-2. This may be even before the clusters undergo core collapse, based on our simulated cluster that has only just formed. The fact that the cluster is still forming stars in a central disk is consistent with the idea that the central stellar density of $>5\times 10^{5}$ $\mathrm{M}_{\odot}$ pc-2 is high enough to prevent pre-SN stellar feedback from expelling the gas (Grudić et al., 2019).
@@
The number density profile shows a shallow central profile with a power-law index close to $-0.5$ on a scale of $\sim 0.1$ pc that is present throughout the cluster assembly. The spatially resolved scale of the stellar component in the simulation is a few $0.01$ pc or less, however the star sampling routine places stars in a Gaussian kernel of $\sigma=0.1$ pc that may affect the central-most density profile as well. Such cored profiles are, however, also a feature in N-body simulations of repeated star cluster mergers (Antonini et al., 2012; Rantala et al., 2024). The relaxation time of the core is approximately one Myr, therefore star formation and relaxation occur simultaneously, together with mass segregation.
@@
We also show a comparison to the best-fit surface density and V-band surface brightness profiles of the young massive cluster R136. The inner radius is limited to the range of data typically used in the fits such as provided by Mackey & Gilmore (2003) and McLaughlin & van der Marel (2005). As with young clusters in general (McLaughlin & van der Marel, 2005), the radial surface brightness profile of R136 is well fit with an Elson-Fall-Freeman profile (Elson et al., 1987). The profile of R136 has an outer power-law slope of the order of $-2$ within the central $\sim 5$ pc (Mackey & Gilmore, 2003; McLaughlin & van der Marel, 2005) and possibly a small central core of $\sim 0.3$ pc. The central R136 cluster is less massive than our simulated cluster by a factor of a few, nonetheless its surface density and light profiles are qualitatively very similar to our simulated cluster. The central mass density of R136 has been estimated to be between $10^{4}$$10^{7}$ $\mathrm{M}_{\odot}$ pc-3 (Selman & Melnick, 2013), depending on the assumed size of the central core. Such values are supported by our hydrodynamical results throughout the cluster assembly.
@@
Figure 4: The radial velocity distribution of the most massive cluster in 0.1 Myr steps. The lines are coloured with the mean stellar age of the cluster, increasing from light to dark. The panels show the 3D velocity dispersion (a), the velocity anisotropy (b), the rotation velocity along a ±$\pm 1$ pc slit along an inclined plane of rotation (c) and the tangential velocity (d). Each line is a mean of 100 bootstrapped samples oriented in a random inclination. The dashed red line in panel b shows the final anisotropy profile without the removal of the bulk motion, $\beta_{\mathrm{rot}}$. The vertical bars in panel a show the projected $R_{\mathrm{eff}}$.

### v2

@@ 3.1 Properties of young star clusters
Approximately 850 star clusters that fill our selection criteria form during the $\sim 7.5$ Myr of starburst. Star clusters above $\sim 300$ $\mathrm{M}_{\odot}$ exhibit a cluster mass function that has a power-law index between 1.9-1.9 and 2-2 throughout the starburst, fit using 20 bins with equal number of 40–50 clusters per bin. The slope of the mass function agrees with observed mass functions and luminosity functions across environments from low-metallicity galaxies such as the Magellanic Clouds to starburst systems (Elmegreen & Falgarone, 1996; Elmegreen & Efremov, 1997; Zhang & Fall, 1999; Hunter et al., 2003; Larsen, 2009; Randriamanakoto et al., 2013).
By the final snapshot, the majority of the clusters are already gas-free while all clusters have mean stellar ages less than 4.3 Myr. The most massive cluster at a mean stellar age of 1.9 Myr still has a gas-embedded core that is forming stars. While the tail end of the starburst is still ongoing, we limit the analysis only until this epoch. The majority of the stellar mass has already formed, and the main stellar collision product (Section 3.1.2) in the most massive cluster has not yet reached a mass significantly above 1000 $\mathrm{M}_{\odot}$, i.e. more than a factor of two beyond the upper stellar mass limit adopted in our VMS models (see Section 2.1.4). Fig. 1 shows the projected half-mass radii ($R_{\mathrm{eff}}$) and the corresponding surface densities $\Sigma_{\mathrm{eff}}=M_{\mathrm{cluster}}/(2\pi R_{\mathrm{eff}}^{2})$ in the xxyy plane for the simulated gravitationally bound star clusters in the final snapshot. The results are compared to masses and half-light radii of young (<10<10 Myr) star clusters in the LEGUS survey (Brown & Gnedin, 2021) and of resolved or marginally resolved star clusters measured through gravitational lensing at redshifts $z\gtrsim 2.4$ (Vanzella et al. 2022b, a, 2023; Adamo et al. 2024; Mowla et al. 2024; Messa et al. 2025). The surface densities in the observed clusters have been computed using the reported half-light radii.
The majority of the simulated star clusters show compact sizes of $R_{\mathrm{eff}}\sim 0.1$11 pc, though only the most massive cluster reaches surface densities of close to $10^{5}$ $\mathrm{M}_{\odot}$ pc-2. The range of initial sizes is in general agreement with compact sizes found in our previous studies in Lahén et al. (2020a, resolution 4 $\mathrm{M}_{\odot}$), Lahén et al. (2024, fully sampled IMF, softened gravity) and Lahén et al. (2025, fully sampled IMF, incl. collisional dynamics with ketju) for young clusters. Cloud-scale simulations of massive star clusters often also report compact sizes, e.g. in Tsang & Milosavljević (2018); Polak et al. (2024). Clusters in the new simulation that have initial masses of a few $10^{2}$$10^{5}$ $\mathrm{M}_{\odot}$ are more extended on average by a factor of 1.5–2 compared to clusters of similar mass and age in the previous setup of Lahén et al. (2024) without collisional dynamics near massive stars. This difference is also seen as a slightly steeper mass-size relation slope of 0.32 compared to 0.26 fit to the clusters in Lahén et al. (2024). Similar to the results of Lahén et al. (2025) for low-mass clusters, the fully exposed clusters undergo rapid size evolution due to internal gravitational dynamics already during the very first Myrs.
@@ 3.1.1 Stellar collisions and two TDEs
In total $\sim 12500$ stars more massive than 88 $\mathrm{M}_{\odot}$ form throughout the starburst. 69 are more massive than $100$ $\mathrm{M}_{\odot}$. When stellar collisions occur in our simulation, only one star grows per cluster after segregating into the cluster core, similar to results in the literature (Portegies Zwart et al., 1999; Baumgardt & Klessen, 2011; Fujii & Portegies Zwart, 2013b; Rantala et al., 2024). In two cases the merger product of two stars below the limit of 100 $\mathrm{M}_{\odot}$ exceeds the VMS limit. The rest of the events that grow VMSs start with a star originally above 100 $\mathrm{M}_{\odot}$. The most massive star cluster hosts >2000>2000 stars more massive than 88 $\mathrm{M}_{\odot}$, up to $\sim 18$ of which are more massive than $100$ $\mathrm{M}_{\odot}$ and present in the cluster at the same time. 12 remain in the cluster in the final snapshot. 13 of the VMSs form after the main progenitor has reached 25% of its final mass and they form within $\sim 1$ pc of the cluster centre. The rest arrive via sub-cluster mergers. By the last snapshot, $78$ stars in the entire galaxy more massive than $40$ $\mathrm{M}_{\odot}$ have collapsed as black holes, out of which $8$ remain in the most massive cluster. 20 stars have exploded as pair-instability supernovae.
@@
We track the collisions of stars throughout the simulation for pairs that include a star or a black hole that had an initial stellar mass of at least 99 $\mathrm{M}_{\odot}$. During the $\sim 7$ Myr of starburst there are in total $90$ collisions throughout the simulation, out of which $16$ involve two components both of which are more massive than $40$ $\mathrm{M}_{\odot}$. Almost $30\%$ of the collisions are between fairly equal mass objects (mass ratio >1/3>1/3). We note that compared to the hierarchical direct N-body simulations of Rantala et al. (2024), we do not see the growth of VMSs through collisions in the in-falling sub-clusters during the assembly of the most massive cluster. This may be due to somewhat delayed mass segregation and sub-cluster core collapse in our collisional dynamics implementation, compared to full direct N-body (see Fig. 1 in Lahén et al. 2025). The core collapse time scale, for instance, affects when VMS growth can occur in the sub-clusters (Fujii & Portegies Zwart, 2013b). Optionally the spread in stellar ages in each cluster due to the star formation history may delay the birth of VMSs and their expansion into the giant phase compared to the uniform age initial conditions adopted in N-body models. This shifts the collisional growth of VMSs to the later, more centrally concentrated stage of cluster assembly.
@@ 3.1.2 The most massive star
In the most massive cluster, one star grows to a peak mass of 1046 $\mathrm{M}_{\odot}$ through repeated collisions with other massive stars. It began its life as a 150 $\mathrm{M}_{\odot}$ star, and consequently merged with stars of 115 $\mathrm{M}_{\odot}$, 170 $\mathrm{M}_{\odot}$, 191 $\mathrm{M}_{\odot}$, 232 $\mathrm{M}_{\odot}$, 153 $\mathrm{M}_{\odot}$ and 44 $\mathrm{M}_{\odot}$ during a period of 1 Myr. The collisions are spaced by 0.1 Myr, 0.15 Myr, 0.4 Myr, 0.3 Myr and 4.4 kyr. The most massive star also merged with 13 low-mass stars (<0.4<0.4 $\mathrm{M}_{\odot}$), increasingly frequently during the final 0.2 Myr as it grew to the supergiant phase. We limit the analysis to this epoch, however the star would still have the potential to collide with a variety of other massive stars, including the most massive normal (i.e. not collisionally formed) VMS of 330 $\mathrm{M}_{\odot}$. This star is at a distance of only 0.05 pc from the cluster core and the most massive star in the final snapshot.
@@
The final VMS mass that far exceeds the upper stellar mass limit of the adopted IMF, and its growth being dominated by collisions with other VMSs, are consistent with hierarchical direct N-body models of star cluster assembly (Rantala et al., 2024). All of the individual collision partners are within the range of estimated initial masses of the most massive stars in R136, therefore the physical scenario for the VMS growth through collisions is very well motivated based on observations. Why no such massive stars that could be collision products are observed in R136 may be due to its too low central/effective density, or its higher metallicity (factor of tens) that allows stronger mass-loss and smaller stellar radii at evolved stages.
@@
Our resulting growth of the most massive star is less extreme than what is found in idealised cloud-scale hydrodynamical simulations (e.g. Reinoso et al. 2023; Fujii et al. 2024; Chon & Omukai 2025). Such simulations of dense star-forming clouds find VMSs in excess of $10^{3}$$10^{4}$ $\mathrm{M}_{\odot}$ forming through collisions and/or gas accretion. Future work with improved treatment for the evolution of the most massive stars will help address whether this difference is a result of the galactic environment and more complex cloud structure in our model, the stellar evolution models, the adopted IMF, or perhaps something else.
@@
Though we have refrained from making assumptions about the detailed evolution of stars beyond $500$ $\mathrm{M}_{\odot}$, an order of magnitude estimate for the final mass of the most massive star and its black hole remnant can be made based on simplified arguments about the mass and metallicity dependent wind mass loss e.g. by Vink (2018) and Sabhahit et al. (2023). The most massive star undergoes its last major collision at an age of $\sim 1.2$ Myr, and it can be expected to still live for one million years. A $\sim 1000$ $\mathrm{M}_{\odot}$ star at $\sim 0.01$ $\mathrm{Z}_{\odot}$ would initially lose mass in optically thick winds at a rate between $\sim 10^{-4.3}$$10^{-3.7}$ $\mathrm{M}_{\odot}$ yr-1 depending on the rate estimate. The star would lose in total between $\sim 25$$160$ $\mathrm{M}_{\odot}$ during the final 1 Myr of its life, resulting in a black hole with a mass of at least $\sim 840$ $\mathrm{M}_{\odot}$. Even if the star would be rotating fast as a consequence of its merger history and therefore lose mass at a factor of 2–3 faster rate as argued by Gieles et al. (2025), it would still result in a final black hole mass of at least $640$ $\mathrm{M}_{\odot}$. Additionally, if the stellar collisions were implemented with a prescription for mass loss, we can approximate the star to end up with 10% lower maximum mass, $\sim 900$ $\mathrm{M}_{\odot}$, after the last major collision. This would further decrease the resulting black hole mass to a minimum of 600 $\mathrm{M}_{\odot}$ (rotating VMS with Sabhahit et al. 2023 mass loss rate) or 770 $\mathrm{M}_{\odot}$ (without rotation). The result would therefore be an intermediate mass black hole in a proto-globular cluster-like system. Whether such a black hole would be retained in the cluster or the galaxy over several Gyrs would require a more detailed modelling of black hole–black hole interactions and mergers in the simulation (see previous studies e.g. by Amaro-Seoane et al. 2014, Arca Sedda et al. 2021, Rantala et al. 2024).
@@
Figure 3: The radial projected profiles of stellar mass surface density (a), V-band surface brightness (b), 3D mass density (c) and stellar number density (d) of the most massive cluster at the four epochs shown in Fig. 2 and in the last snapshot when the cluster has a mean stellar age of 1.9 Myr. The solid lines in panels a, c and d indicate the gravitationally bound stellar component and the dashed lines show the total stellar mass in the region. In panel b the solid and dashed lines show the attenuated and the unattenuated surface brightness profiles, respectively. The power-law slopes of 0.5-0.5, 1-1, 2-2 and 3-3 are indicated in the panels. The best-fit surface density and brightness profiles of R136 from McLaughlin & van der Marel (2005) are shown in black in panels a and b, and the right-side y-axis in panel b shows the surface brightness scale in mag arcsec-2. The vertical bars at the bottom of the panels show the projected half-mass radii of the bound component (panels a, c and d) and the projected half-light radii of the attenuated and unattenuated light profiles (b) with respective linestyles.
@@ 3.2 Properties of the most massive cluster
The stellar light of the main cluster is attenuated and the other massive clump in the images remains the brightest cluster in V-band for the duration of this coalescence, irrespective of the masking applied in the centring. The V-band attenuation $A_{V}$ is $\sim 0.9$ mag at these epochs, computed as the AB-magnitude difference between the unattenuated and attenuated images within a radius of 1 pc of the centre of light of the unattenuated image. This is a moderate value for observed star clusters and not as extreme as the most attenuated clusters observed in some dwarf galaxies (e.g. Calzetti et al. 2023). The reddening E(BV)E(B-V) within the same 1 pc region is 0.7 and 0.5 in the two first epochs, similar to what is typically found for young clusters in dwarf galaxies or starbursting systems (E(BV)E(B-V) up to 0.7 e.g. in Adamo et al. 2020; Whitmore et al. 2020).
The last two epochs in Fig. 2 show centrally concentrated star formation. The central 1 pc has a star-forming disky gas component until the last snapshot. The gas disk is being fed along a gaseous filament that extends toward the cluster from north-west in the rightmost panels. $A_{V}$ peaks at $\sim 1.4$ during the last epochs, and E(BV)E(B-V) drops to 0.3–0.4.
@@ 3.2.1 Density, surface density and surface brightness profiles
The density and brightness profiles of the most massive cluster are consistent with power-law shapes especially in the outer parts and at later epochs when the stellar distribution is less clumpy (see Fig. 2). The infalling clusters in our simulation can be seen as an increase in density at intermediate radii (1133 pc) at earlier epochs. The outer radial power-law slope in surface density is 2.2-2.2, as expected from models of hierarchical sub-clusters assembly that predict slopes that approach 2-2 (Sills et al., 2018; Grudić et al., 2018; Rantala et al., 2024). The outer power-law has a respective density slope of slightly shallower than 3-3 and continues well into the sub-parsec regime. The final central 3D mass density, 3D number density and surface mass density (inner $\sim 1000$ stars) are a factor of $\sim 190$, 10 and 10 times higher than the half-mass values shown in Fig. 1, respectively. The unattenuated central surface brightness profile likewise has a steep inner slope, dominated by the most massive stars in the cluster core. Based on these simulated brightness and density profiles, we therefore infer that the central surface densities $\Sigma_{0}$ of the observed high-redshift clusters may be at least an order of magnitude higher than the effective values $\Sigma_{\mathrm{eff}}$, i.e. $\Sigma_{0}>10^{6}$ $\mathrm{M}_{\odot}$ pc-2. This may be even before the clusters undergo core collapse, based on our simulated cluster that has only just formed. The fact that the cluster is still forming stars in a central disk is consistent with the idea that the central stellar density of $5\times 10^{5}$ $\mathrm{M}_{\odot}$ pc-2 is high enough to prevent pre-SN stellar feedback from expelling the gas (Grudić et al., 2019).
@@
The number density profile shows a shallow central profile that is an expected feature in N-body simulations of repeated star cluster mergers (Antonini et al., 2012; Rantala et al., 2024). The relaxation time of the core is approximately one Myr, therefore star formation and relaxation occur simultaneously, together with mass segregation. The spatially resolved scale of the stellar component in the simulation is a few $0.01$ pc or less, however we note that the star sampling routine places stars in a Gaussian kernel of $\sigma=0.1$ pc that impacts the central-most density profile of the youngest stars. We therefore refrain from making quantitative statements regarding the density profile in the core region.
@@
We also show a comparison to the best-fit surface density and V-band surface brightness profiles of the young massive cluster R136. The inner radius is limited to the range of data typically used in the fits such as provided by Mackey & Gilmore (2003) and McLaughlin & van der Marel (2005). As with young clusters in general (McLaughlin & van der Marel, 2005), the radial surface brightness profile of R136 is well fit with an Elson-Fall-Freeman profile (Elson et al., 1987). The profile of R136 has an outer power-law slope of the order of 2-2 within the central $\sim 5$ pc (Mackey & Gilmore, 2003; McLaughlin & van der Marel, 2005) and possibly a small central core of $\sim 0.3$ pc. The central R136 cluster is less massive than our simulated cluster by a factor of a few, nonetheless its surface density and light profiles are qualitatively very similar to our simulated cluster. The central mass density of R136 has been estimated to be between $10^{4}$$10^{7}$ $\mathrm{M}_{\odot}$ pc-3 (Selman & Melnick, 2013), depending on the assumed size of the central core. Such values are supported by our hydrodynamical results throughout the cluster assembly.
@@
Figure 4: The radial velocity distribution of the most massive cluster in 0.1 Myr steps. The lines are coloured with the mean stellar age of the cluster, increasing from light to dark. The panels show the 3D velocity dispersion (a), the velocity anisotropy (b), the rotation velocity along a $\pm 1$ pc slit along an inclined plane of rotation (c) and the tangential velocity (d). Each line is a mean of 100 bootstrapped samples oriented in a random inclination. The dashed line in panel b shows the final anisotropy profile without the removal of the bulk motion, $\beta_{\mathrm{rot}}$. The vertical bars in panel a show the projected $R_{\mathrm{eff}}$.